# H-SPLID: HSIC-based Saliency Preserving Latent Information Decomposition

### v1

@@ Abstract
We introduce H-SPLID, a novel algorithm for learning salient feature representations through the explicit decomposition of salient and non-salient features into separate spaces. We show that H-SPLID promotes learning low-dimensional, task-relevant features. We prove that the expected prediction deviation under input perturbations is upper-bounded by the dimension of the salient subspace and the Hilbert-Schmidt Independence Criterion (HSIC) between inputs and representations. This establishes a link between robustness and latent representation compression in terms of the dimensionality and information preserved. Empirical evaluations on image classification tasks show that models trained with H-SPLID primarily rely on salient input components, as indicated by reduced sensitivity to perturbations affecting non-salient features, such as image backgrounds. Our code is available at https://github.com/neu-spiral/H-SPLID.
@@ 1 Introduction
The acquisition of salient, task-relevant features from high-dimensional inputs constitutes a fundamental challenge in representation learning. Such features offer multiple advantages, including reduced dimensionality [2], enhanced generalization and transferability [42, 30], and improved robustness [22, 12]. Nevertheless, learning true salient features remains challenging, as many neural networks operate within a single, entangled latent space that mixes task-relevant signals with redundant information [6, 38]. We illustrate this sensitivity using a simple diagnostic test in Figure 1: a classifier trained to predict the left digit in an image of double digits should ignore perturbations to the right digit, which is irrelevant to the label. However, in practice, we observe that neural networks with high test classification accuracy on the left digits exhibit a significant performance drop when subjected to a high-magnitude adversarial PGD [34] attack ($\epsilon=1.0$) on the right digits, revealing their dependence on irrelevant, non-salient features. This corroborates several empirical and theoretical studies [3, 36, 19] showing that redundant dimensions enhance vulnerability to attacks. Driven by these findings, we introduce H-SPLID (HSIC-based Saliency-Preserving Latent Information Decomposition), a new method that learns salient features by explicitly decomposing the latent space into two subspaces coupled with information compression regularization: a low-dimensional salient space, which contains features essential for classification, and a non-salient space, which captures the remaining input variability. Training the same neural network as above with H-SPLID alleviates the dependence on irrelevant features to a large degree, as shown on the right side of Figure 1. Crucially, H-SPLID is significantly less sensitive to right digit perturbations, without any prior knowledge of the redundant region or adversarial training.
@@ 3.3 Regularizing and Preserving the Separated Features
subject to $\mathbf{M}_{n}=I-\mathbf{M}_{s},\quad\mathbf{z}_{i}=f_{\boldsymbol{\psi}}(\mathbf{x}_{i}),~\forall i\in\{1,\dots,m\}.$ (6b)
@@ 3.4 The H-SPLID Algorithm
${\boldsymbol{\theta}}^{(t)}\leftarrow{\boldsymbol{\theta}}^{t-1}-\eta\nabla_{{\boldsymbol{\theta}}}\mathcal{L}_{\text{ST}}(\mathcal{B};{\boldsymbol{\theta}}^{(t-1)},\mathbf{M}_{s}^{(t-1)},\mathbf{M}_{n}^{(t-1)})$
@@ 3.5 Theoretical Guarantees
Let $\mathbf{x}$ be sampled from a tMVN distribution $\mathbf{x}\sim\mathcal{N}_{R}(0,\sigma^{2}I_{d})$ with radius $R$, and let the neural network $h_{\boldsymbol{\theta}}:\mathbb{R}^{d}\to\mathbb{R}^{k}$ be differentiable almost everywhere with an $L$-Lipschitz encoder $f_{\psi}$ and a bounded linear output layer with $\|\mathbf{W}\|_{\infty}\leq B$. Suppose the RKHSs associated with $K_{x}$ and $K_{z}$ satisfy Assumption 3.1, with kernel sup-norm bounds $K_{\mathcal{F}},K_{\mathcal{G}}$. Then, given the count of non-zero entries of the salient mask $s\coloneqq\|\mathbf{M}_{s}\|_{0}$, for any perturbation $\delta$ with $\|\delta\|_{2}\leq r$, the expected output deviation satisfies
$\mathbb{E}_{\mathbf{x}}\left[\|h_{\boldsymbol{\theta}}(\mathbf{x}+\delta)-h_{\boldsymbol{\theta}}(\mathbf{x})\|_{2}\right]\leq\frac{rRB\sqrt{ks}(LR+\|f_{\boldsymbol{\psi}}(0)\|_{2})}{\sigma^{2}K_{\mathcal{F}}K_{\mathcal{G}}}\cdot\operatorname{HSIC}(\mathbf{x},\mathbf{z}_{s})+o(r),$ (9)
@@
The proof of Theorem˜3.2 can be found in Appendix˜C. Intuitively, with stronger information compression imposed by minimizing both the HSIC term (thereby reducing $\operatorname{HSIC}(\mathbf{x},\mathbf{z}_{s}))$ and the masks (thereby reducing the salient mask support $s\coloneqq\|\mathbf{M}_{s}\|_{0}$), the model is forced to rely only on salient features: Theorem˜3.2 suggests that, in this case, the perturbation is more likely to end up in the non-salient space, and the majority of the attack does not contribute to the change of the output of the neural network, as $\|h_{\boldsymbol{\theta}}(\mathbf{x}+\delta)-h_{\boldsymbol{\theta}}(\mathbf{x})\|$ stays small for any perturbation $\delta$ with $\|\delta\|_{2}\leq r$.
From a technical perspective, our theorem differs from Wang et al. [56] in two aspects. First, we sharpen the dependence of the upper bound on the power of the perturbation; this allows us to explicitly link it to the dimension of the salient mask $s$. Second, we extend their binary classification framework (i.e., $k=1$ with one output value) to multi-class classification (with arbitrary $k$) to cover a wider range of classification models. Additionally, we can quantify how perturbations will trigger prediction changes from the volume of the entire input domain. In particular, we define the salient-active region as $\mathcal{X}_{s}(\epsilon):=\left\{\mathbf{x}\in\mathcal{X}_{R}:\|\nabla_{\mathbf{x}}h_{\boldsymbol{\theta}}(\mathbf{x})\|_{F}>\epsilon\right\}.$ Under the tMVN distribution, the probability that an input falls into this region equals its measure: $\mu\left(\mathcal{X}_{s}(\epsilon)\right)=\mathbb{P}\left(\|\nabla_{\mathbf{x}}h_{\boldsymbol{\theta}}(\mathbf{x})\|_{F}>\epsilon\right).$ We can bound this probability as follows:
@@ Appendix B H-SPLID Pseudocode
7:   Compute loss $\mathcal{L}_{\text{ST}}(\mathcal{B};{\boldsymbol{\theta}},\mathbf{M}_{s},I-\mathbf{M}_{s})$
8:   Update ${\boldsymbol{\theta}}\leftarrow{\boldsymbol{\theta}}-\eta\nabla_{{\boldsymbol{\theta}}}\mathcal{L}_{\text{ST}}$
@@
Using first-order Taylor expansion with sufficiently small perturbation $\|\delta\|_{2}\leq r$, we have
@@
Thus, by combining eq.˜31 and eq.˜32, it follows
$\mathbb{E}\left[\|h_{\boldsymbol{\theta}}(X+\delta)-h_{\boldsymbol{\theta}}(\mathbf{x})\|_{2}\right]\leq$ $r\cdot\mathbb{E}\left[\|J_{h_{\boldsymbol{\theta}}}(\mathbf{x})\|_{F}\right]$ (33)
$\leq$ $\frac{rRB\sqrt{ks}(LR+\|f_{\boldsymbol{\psi}}(0)\|_{2})}{\sigma^{2}K_{\mathcal{F}}K_{\mathcal{G}}}\cdot\operatorname{HSIC}(\mathbf{x},\mathbf{z}_{s})+o(r),$ (34)
where in the last step we replace the term $N_{\mathcal{X}},N_{\mathcal{Z}}$ and use the fact $\frac{\epsilon_{R}}{\sigma^{2}}=o(1)$ due to the exponentially decaying term.
@@
Thus, plugging eq.˜62 to eq.˜63, we have

### v2

@@ Abstract
We introduce H-SPLID, a novel algorithm for learning salient feature representations through the explicit decomposition of salient and non-salient features into separate spaces. We show that H-SPLID promotes learning low-dimensional, task-relevant features. We prove that the expected prediction deviation under input perturbations is upper-bounded by the dimension of the salient subspace and the Hilbert-Schmidt Independence Criterion (HSIC) between inputs and representations. This establishes a link between robustness and latent representation compression in terms of the dimensionality and information preserved. Empirical evaluations on image classification tasks show that models trained with H-SPLID primarily rely on salient input components, as indicated by reduced sensitivity to perturbations affecting non-salient features, such as image backgrounds.
@@ 1 Introduction
The acquisition of salient, task-relevant features from high-dimensional inputs constitutes a fundamental challenge in representation learning. Such features offer multiple advantages, including reduced dimensionality [1], enhanced generalization and transferability [42, 30], and improved robustness [22, 12]. Nevertheless, learning true salient features remains challenging, as many neural networks operate within a single, entangled latent space that mixes task-relevant signals with redundant information [6, 38]. We illustrate this sensitivity using a simple diagnostic test in Figure 1: a classifier trained to predict the left digit in an image of double digits should ignore perturbations to the right digit, which is irrelevant to the label. However, in practice, we observe that neural networks with high test classification accuracy on the left digits exhibit a significant performance drop when subjected to a high-magnitude adversarial PGD [34] attack ($\epsilon=1.0$) on the right digits, revealing their dependence on irrelevant, non-salient features. This corroborates several empirical and theoretical studies [3, 36, 19] showing that redundant dimensions enhance vulnerability to attacks. Driven by these findings, we introduce H-SPLID (HSIC-based Saliency-Preserving Latent Information Decomposition), a new method that learns salient features by explicitly decomposing the latent space into two subspaces coupled with information compression regularization: a low-dimensional salient space, which contains features essential for classification, and a non-salient space, which captures the remaining input variability.
Training the same neural network as above with H-SPLID alleviates the dependence on irrelevant features to a large degree, as shown on the right side of Figure 1. Crucially, H-SPLID is significantly less sensitive to right digit perturbations, without any prior knowledge of the redundant region or adversarial training.
@@ 3.3 Regularizing and Preserving the Separated Features
subject to $\mathbf{M}_{n}=I-\mathbf{M}_{s},\quad\mathbf{z}_{i}=f_{\boldsymbol{\psi}}(\mathbf{x}_{i}),~\forall i\in\{1,\dots,n\}.$ (6b)
@@ 3.4 The H-SPLID Algorithm
${\boldsymbol{\theta}}^{(t)}\leftarrow{\boldsymbol{\theta}}^{t-1}-\eta\nabla_{{\boldsymbol{\theta}}}\mathcal{L}(\mathcal{B};{\boldsymbol{\theta}}^{(t-1)},\mathbf{M}_{s}^{(t-1)},\mathbf{M}_{n}^{(t-1)})$
@@ 3.5 Theoretical Guarantees
Let $\mathbf{x}$ be sampled from a tMVN distribution $\mathbf{x}\sim\mathcal{N}_{R}(0,\sigma^{2}I_{d})$, and let the neural network $h_{\boldsymbol{\theta}}:\mathbb{R}^{d}\to\mathbb{R}^{k}$ be differentiable almost everywhere with an $L$-Lipschitz encoder $f_{\psi}$ and a bounded linear output layer with $\|\mathbf{W}\|_{\infty}\leq B$. Suppose the RKHSs associated with $K_{x}$ and $K_{z}$ satisfy Assumption 3.1, with kernel sup-norm bounds $K_{\mathcal{F}},K_{\mathcal{G}}$. Let $s\coloneqq\|\mathbf{M}_{s}\|_{0}$ be the count of non-zero entries of the salient mask. Then, for all perturbation maps $\delta:\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that $\|\delta(\mathbf{x})\|_{2}\leq r$ for all $\mathbf{x}\in\mathcal{X}$}, the expected output deviation satisfies
$\mathbb{E}_{\mathbf{x}}\left[\|h_{\boldsymbol{\theta}}(\mathbf{x}+\delta(\mathbf{x}))-h_{\boldsymbol{\theta}}(\mathbf{x})\|_{2}\right]\leq\frac{rRB\sqrt{ks}(LR+\|f_{\boldsymbol{\psi}}(0)\|_{2})}{\sigma^{2}K_{\mathcal{F}}K_{\mathcal{G}}}\cdot\operatorname{HSIC}(\mathbf{x},\mathbf{z}_{s})+o(r),$ (9)
@@
The proof of Theorem˜3.2 can be found in Appendix˜C. Intuitively, with stronger information compression imposed by minimizing both the HSIC term (thereby reducing $\operatorname{HSIC}(\mathbf{x},\mathbf{z}_{s}))$ and the masks (thereby reducing the salient mask support $s\coloneqq\|\mathbf{M}_{s}\|_{0}$), the model is forced to rely only on salient features: Theorem˜3.2 suggests that, in this case, the perturbation is more likely to end up in the non-salient space, and the majority of the attack does not contribute to the change of the output of the neural network, as $\|h_{\boldsymbol{\theta}}(\mathbf{x}+\delta(\mathbf{x}))-h_{\boldsymbol{\theta}}(\mathbf{x})\|$ stays small for any perturbation map $\delta(\cdot)$ with $\|\delta(\mathbf{x})\|_{2}\leq r$. From a technical perspective, our theorem differs from Wang et al. [56] in two aspects. First, we sharpen the dependence of the upper bound on the power of the perturbation; this allows us to explicitly link it to the dimension of the salient mask $s$. Second, we extend their binary classification framework (i.e., $k=1$ with one output value) to multi-class classification (with arbitrary $k$) to cover a wider range of classification models.
Additionally, we can quantify how perturbations will trigger prediction changes from the volume of the entire input domain. In particular, we define the salient-active region as $\mathcal{X}_{s}(\epsilon):=\left\{\mathbf{x}\in\mathcal{X}_{R}:\|\nabla_{\mathbf{x}}h_{\boldsymbol{\theta}}(\mathbf{x})\|_{F}>\epsilon\right\}.$ Under the tMVN distribution, the probability that an input falls into this region equals its measure: $\mu\left(\mathcal{X}_{s}(\epsilon)\right)=\mathbb{P}\left(\|\nabla_{\mathbf{x}}h_{\boldsymbol{\theta}}(\mathbf{x})\|_{F}>\epsilon\right).$ We can bound this probability as follows:
@@ Appendix B H-SPLID Pseudocode
7:   Compute loss $\mathcal{L}(\mathcal{B};{\boldsymbol{\theta}},\mathbf{M}_{s},I-\mathbf{M}_{s})$
8:   Update ${\boldsymbol{\theta}}\leftarrow{\boldsymbol{\theta}}-\eta\nabla_{{\boldsymbol{\theta}}}\mathcal{L}$
@@
Consider the first-order Taylor expansion in a Euclidian ball of radius $r$ around $\mathbf{x}\in\mathcal{X}$: that is,
@@
for all $\mathbf{x}\in\mathcal{X}$ and all $\delta\in\mathbb{R}^{d}$ s.t. $\|\delta\|_{2}\leq r$.
Consider now a measurable perturbation function on $\mathcal{X}$ as $\delta:\mathbb{R}^{d}\to\mathbb{R}^{d}$, such that
$\sup_{\mathbf{x}\in\mathcal{X}}\|\delta(\mathbf{x})\|_{2}\;\leq\;r.$ (33)
As Eq. (32) holds for all $\mathbf{x},\delta$ pairs, for all $\mathbf{x}\in\mathcal{X}$, we have that:
$h_{\boldsymbol{\theta}}(\mathbf{x}+\delta(\mathbf{x}))-h_{\boldsymbol{\theta}}(\mathbf{x})=J_{h_{\boldsymbol{\theta}}}(\mathbf{x})\delta(\mathbf{x})+o(r).$ (34)
We thus have that, for all $\mathbf{x}\in\mathcal{X}$:
$\big\|h_{\boldsymbol{\theta}}(\mathbf{x}+\delta(\mathbf{x}))-h_{\boldsymbol{\theta}}(\mathbf{x})\big\|_{2}$ $\leq\big\|J_{h_{\boldsymbol{\theta}}}(\mathbf{x})\,\delta(\mathbf{x})\big\|_{2}+o(r)$ (35)
$\leq\big\|J_{h_{\boldsymbol{\theta}}}(\mathbf{x})\big\|_{F}\,\big\|\delta(\mathbf{x})\big\|_{2}+o(r)$ $\text{by Cauchy-Schwartz},$ (36)
[…]
Hence, it follows that:
$\mathbb{E}\left[\big\|h_{\boldsymbol{\theta}}(X+\delta(X))-h_{\boldsymbol{\theta}}(X)\big\|_{2}\right]$ $\leq r\cdot\mathbb{E}\left[\big\|J_{h_{\boldsymbol{\theta}}}(X)\big\|_{F}\right]+o(r)$ (38)
$\leq\frac{rRB\sqrt{ks}(LR+\|f_{\boldsymbol{\psi}}(0)\|_{2})}{\sigma^{2}K_{\mathcal{F}}K_{\mathcal{G}}}\cdot\operatorname{HSIC}(X,\mathbf{z}_{s})+o(r),$ (39)
where the last inequality follows from Eq.˜31 and the fact $\frac{\epsilon_{R}}{\sigma^{2}}=o(1)$, due to the exponentially decaying term. ∎
@@
Thus, plugging eq.˜67 to eq.˜68, we have